OpenStudy (anonymous):

What is the x value in the proportion x/x-6 = 4/7? This is really confusing me since I cross multiplied

OpenStudy (diyadiya):

Yes! cross multiply what do you get?

OpenStudy (anonymous):

-24 and 7x

OpenStudy (diyadiya):

\[\frac{x}{x-6}= \frac{4}{7}\]\[7x = 4 (x-6)\]

OpenStudy (diyadiya):

\[7x = 4x - 24\] Is this clear?

OpenStudy (anonymous):

that's what i did wrong. i multiplied x-6 and 4 the wrong way.

OpenStudy (diyadiya):

Yeah! hope you can continue from here.. can you?

OpenStudy (diyadiya):

we distribute 4(x-6) like a(b-c) = ab -ac